# Step-by-Step Reasoning to Solve Grid Puzzles: Where do LLMs Falter?

### v1

@@ 1 Introduction
Recent advancements in LLMs such as GPT-4, Gemini, Claude-3 Anthropic (2024), Llama-2 Touvron et al. (2023), and Mistral Jiang et al. (2023) have achieved remarkable performance on a wide range of Natural Language Understanding (NLU) tasks previously thought to be exclusive to human. Beyond NLU, exploring LLMs’ logical reasoning abilities Liu et al. (2021); Saparov and He (2022); Parmar et al. (2024); Patel et al. (2024) on complex tasks such as puzzle-solving is under-explored. Past attempts have been made to evaluate models on logic-intensive grid-based puzzle-solving. However, they either do not focus on evaluating LLMs Mitra and Baral (2015); Jabrayilzade and Tekir (2020) or do not evaluate LLMs independently, but rather use neuro-symbolic approaches Ishay et al. (2023) that use external specialized solvers on LLM outputs. Here, we aim to evaluate the puzzle-solving abilities of LLMs by themselves, without the use of any external logic solvers.
@@
To investigate reasoning chains, we manually analyze them (Figure 1) to find fine-grained errors (further details in section 3.3). Based on this, we propose a new error taxonomy comprising five broad categories, and nine fine-grained sub-categories (Tables 1 and 2), providing deeper insights into the primary causes of the LLMs’ reasoning failures. However, scaling manual analysis to a larger set is time-consuming and laborious. Hence, we propose to leverage LLMs as auto-evaluators by creating prompts that utilize error taxonomy to automate the analysis of reasoning chains and help in identifying errors (Figure 1). While evaluating w.r.t. manual annotation, our auto-evaluator model achieves $\sim 86\%$ agreement, hence providing quality error categorization.
@@ 2 Related Work
Puzzle-solving task provides detailed insights into LLMs’ logical reasoning. Giadikiaroglou et al. (2024) categorize puzzles into (1) rule-based and (2) rule-less puzzles. Rule-less puzzles include riddles Lin et al. (2021), MCQs Zhao and Anderson (2023), programming puzzles Schuster et al. (2021), and commonsense reasoning puzzles Gu et al. (2023); however, in our work we focus on rule-based puzzles. In rule-based puzzles, past attempts have explored Sudoku Noever and Burdick (2021), Rubik’s Cube, 8-puzzle, Game of 24 Yao et al. (2024), crosswords Yao et al. (2024), chess puzzles Feng et al. (2024), card games Gupta (2023), and BoardgameQA Kazemi et al. (2024). However, grid-based puzzle solving is under-explored. Mitra and Baral (2015) proposed a grid-based puzzle dataset, but this work does not provide any insights into the performance of recent LLMs. Motivated by this, we propose a systematically curated grid-based puzzle dataset, GridPuzzle, and also provide a detailed evaluation of various LLMs in puzzle-solving.
@@ 3.2 Manual Evaluation
To explore where exactly these LLMs falter in performing reasoning, we conduct a detailed manual analysis of the reasoning chains generated by them while solving grid-based puzzles. Our manual analysis process consists of three steps. First, we begin by segmenting the reasoning chains into individual sentences, allowing us to categorize errors more precisely. Second, we identify the premise and conclusion for each sentence and determine their respective correctness. We refrain from subdividing sentences into multiple premises or conclusions to maintain simplicity for finding errors. At last, each sentence is categorized as either containing a single premise and conclusion or being a declarative statement without a conclusion. Then, we begin assessing potential issues or errors in the reasoning chains. Now, we follow an exhaustive approach to create fine-grained error categories. We begin with 30 reasoning chains (6 puzzles x 5 reasoning chains from LLMs) to manually identify potential errors. Next, we categorize these errors in a structured format. We then add another 30 reasoning chains to see if any new types of errors emerge. If new errors are identified, we refine our categories accordingly. This process is repeated until we evaluate a total of 150 reasoning chains and no new types of errors are found. Based on this method, we have carefully filtered and categorized several errors made by LLMs, presenting them as five broad categories and nine sub-categories.
@@ 4.1 Experiments
We evaluate a range of closed-source LLMs including GPT-4-Turbo, Claude-3-Opus, and Gemini-Pro, and open-source models Llama-2-13B, and Mistral-7B-Instruct on GridPuzzle in the Zero-shot-CoT setting Kojima et al. (2022). Our GridPuzzle dataset consists of a set of instances denoted as $\mathcal{P}={<p_{n}^{i\times j},a_{n}>}$, where $p_{n}^{i\times j}$ is $n^{th}$ puzzle instance with grid size of $i\times j$ and $a_{n}$ as a gold answer. We prompt each LLM to generate a reasoning chain before predicting answer $\hat{a}$. To evaluate each model in the Zero-shot-CoT setting, we provide $<I,p_{n}^{i\times j}>$ as input to the model and predict an answer $\hat{a}$ where $I$ is a natural language instruction. The evaluation is conducted on the OpenAI, Google, and Anthropic model versions released in April 2024 with temperature setting 0 for deterministic predictions. NVIDIA A100 GPUs are used for conducting the inference of open-source models with a batch size of 4. The example prompts used for these experiments are provided in Appendix A.
Figure 2: The process of calculating PuzzleEval metrics is described above. The reasoning chains are produced by our 5 LLM models and the gold solution is taken from our GridPuzzle dataset.
@@ 4.2 Metrics
Figure 3: Performance of 5 different LLMs in terms of accuracy on the GridPuzzle dataset.
@@ 5.2 Reasoning Chain Evaluation
Figure 4: The percentage distribution of the broad error categories across the combined reasoning steps of all 5 LLMs. The total number of steps generated by each model is provided inside the round brackets below the model names.
@@
Why is the accuracy low despite the reasoning chains being mostly error-free?
The disparity between accuracy and PuzzleEval arises from the relative location of errors within the reasoning chains. It has been observed that “RR” category reasoning steps mainly occur in the initial half of the chain, leading to a high overall PuzzleEval score. Conversely, errors in the “RW”, “WR”, and “WW” categories typically occur in the latter half of the chain, resulting in incorrect final answers and lower accuracy scores. Based on our error taxonomy, the “RW”, “WR”, and “WW” broad error categories have been further dissected into $6\times 3$ error sub-categories, with their distribution across reasoning steps shown in Figure 5.
Dominant error sub-categories and their correlation with the broad categories.
The most common error sub-category across all heatmaps appears to be the “-” category which refers to the absence of errors. All the reasoning steps with “NC” and “RR” classifications fall in this category. To observe the actual overall trend across all 5 LLMS, the top 10 most common error sub-categories have been listed in the last sub-figure of Figure 5. The top categories ‘a’ and ‘c’ refer to the Wrong Reasoning and the Wrong Elimination sub-categories under the “RW” category. These errors arise when the premise is correct but the LLMs fail to make accurate deductions from it. A number of the top 10 sub-error categories (‘1a’, ‘4a’, ‘4b’, ‘4c’, ‘5a’, ‘5c’, and ‘6a’) emerge from the “WW” category.
@@
Table 5: The results for accuracy and PuzzleEval using GPT-4-Turbo, with and without mitigation strategies for the 30 samples of $3\times 4$ grid-size.
Performance of Open-source vs Proprietary LLMs.
From the results of our objective and subjective metrics, it is evident that the open-source models have lower performance on the grid-puzzle-solving task than the proprietary models. The Llama-2 and Mistral models have the lowest accuracy values and their low performance on the PuzzleEval consistently degrades with the increase in the size and complexity of the grids. The Claude-3, Gemini, and GPT-4 models have higher values of accuracy but their performance across the grid sizes in the PuzzleEval is inconsistent. The disparity in the performance of both kinds of models can be attributed to the difference in their parameter sizes and the low instruction following capabilities of small open-source models.
Discussion on Mitigation Strategy
We conduct a case study on a subset of GridPuzzle, focusing on a 3x4 grid size, utilizing commonly employed prompting techniques to enhance the reasoning capabilities of LLMs. In particular, we use four strategies: (1) Plan-and-Solve Wang et al. (2023), (2) Self-correct Zhang et al. (2024), (3) Self-discover Zhou et al. (2024), and (4) Feedback-Learning. We updated the prompts corresponding to these techniques to include some of our major findings from the reasoning chain evaluations and error categorization analysis as precautionary instructions.
The first strategy Plan-and-Solve prompts the model to first generate a plan to solve the given problem and then follow those steps. The second strategy is inspired by the Self-correct method which uses a combination of self-verification and self-refine to improve reasoning. Next, we used the Self-discover technique which is a 2-step structured reasoning process. Lastly, we created our prompting technique called “Feedback-Learning” by providing specific feedback system instructions to the LLM based on our error taxonomy. The detailed prompt structure is described in Appendix D and the results of these strategies are in Table 5. It is evident from the results that prompting-based strategies are not sufficient to significantly improve the LLM reasoning on the grid-puzzle-solving task. Compared to the rest of the strategies, Self-Discover marginally improves the performance on both accuracy and PuzzleEval. These results indicate the need to develop techniques beyond prompting by having deeper insights from LLMs’ reasoning chains.
@@ 6 Conclusion
We have used AI assistants (Grammarly and ChatGPT) to address the grammatical errors and rephrase the sentences.
@@ Acknowledgement
We extend our gratitude to the Research Computing (RC), and Enterprise Technology at ASU for providing computing resources, and access to the ChatGPT enterprise version for experiments. We acknowledge support by a 2023 Spring Amazon Research Award (ARA).
Appendix A GridPuzzle dataset sample Puzzle
@@
Appendix B GPT-4o Auto Evaluator
@@ Appendix C Evaluation of Reasoning Chains
Appendix D Mitigation Strategy Prompts
@@
The Self-correct strategy merges Self-verify and Self-refine qualities to minimize LLM reasoning errors. It starts with solving the puzzle using a Zero-shot-CoT prompt, followed by prompting the LLM to verify and refine the solution. Finally, it integrates the model’s suggestions with insights from our error taxonomy to enhance the puzzle-solving response. The prompt structure for this strategy is shown in figure 12. The Self-Discover strategy, depicted in figure 13, proved most effective in reducing LLM reasoning errors in puzzle-solving. This approach begins by having the model analyze the problem and potential errors, follows with a list of prescribed reasoning modules, prompts the LLM to select and apply the most suitable module, and concludes by using a structured prompt to solve the puzzle.

### v2

@@ 1 Introduction
Recent advancements in LLMs such as GPT-4, Gemini, Claude-3 Anthropic (2024), Llama-2 Touvron et al. (2023), and Mistral Jiang et al. (2023) have achieved remarkable performance on a wide range of Natural Language Understanding (NLU) tasks previously thought to be exclusive to humans. Beyond NLU, exploring LLMs’ logical reasoning abilities Liu et al. (2021); Saparov and He (2022); Parmar et al. (2024); Patel et al. (2024) on complex tasks such as puzzle-solving is under-explored. Past attempts have been made to evaluate models on logic-intensive grid-based puzzle-solving. However, they either do not focus on evaluating LLMs Mitra and Baral (2015); Jabrayilzade and Tekir (2020) or do not evaluate LLMs independently, but rather use neuro-symbolic approaches Ishay et al. (2023) that use external specialized solvers on LLM outputs. Here, we aim to evaluate the puzzle-solving abilities of LLMs by themselves, without the use of any external logic solvers.
@@
To investigate the reasoning chains, we manually analyze them (Figure 1) to find fine-grained errors (further details in section 3.3). Based on this, we propose a new error taxonomy comprising five broad categories, and nine fine-grained sub-categories (Tables 1 and 2), providing deeper insights into the primary causes of the LLMs’ reasoning failures. However, scaling manual analysis to a larger set is time-consuming and laborious. Hence, we propose to leverage LLMs as auto-evaluators by creating prompts that utilize error taxonomy to automate the analysis of reasoning chains and help in identifying errors (Figure 1). While evaluating w.r.t. manual annotation, our auto-evaluator model achieves $\sim 86\%$ agreement, hence providing quality error categorization.
@@ 2 Related Work
Puzzle-solving task provides insights into LLMs’ logical reasoning. Giadikiaroglou et al. (2024) categorize puzzles into (1) rule-based and (2) rule-less puzzles. Rule-less puzzles include riddles Lin et al. (2021), MCQs Zhao and Anderson (2023), programming puzzles Schuster et al. (2021), and commonsense reasoning puzzles Gu et al. (2023); however, in our work we focus on rule-based puzzles. In rule-based puzzles, past attempts have explored Sudoku Noever and Burdick (2021), Rubik’s Cube, 8-puzzle, Game of 24 Yao et al. (2024), crosswords Yao et al. (2024), chess puzzles Feng et al. (2024), card games Gupta (2023), BoardgameQA Kazemi et al. (2024), and Lateral Thinking Puzzles Huang et al. (2024). However, grid-based puzzle solving is under-explored. Mitra and Baral (2015) proposed a grid-based puzzle dataset and Dziri et al. (2023) studied compositionality in LLMs using Grid Puzzle, but these works do not provide any insights into the performance of recent LLMs. Motivated by this, we propose a systematically curated grid-based puzzle dataset, GridPuzzle, and provide an evaluation of various LLMs in puzzle-solving.
@@ 3.2 Manual Evaluation
To explore where exactly these LLMs falter in performing reasoning, we conduct a detailed manual analysis of the reasoning chains generated by them while solving grid-based puzzles. Details of the annotation guidelines provided to the human evaluators are given in the Appendix D. Our manual analysis process consists of three steps. First, we begin by segmenting the reasoning chains into individual sentences, allowing us to categorize errors more precisely. Second, we identify the premise and conclusion for each sentence and determine their respective correctness. We refrain from subdividing sentences into multiple premises or conclusions to maintain simplicity for finding errors. At last, each sentence is categorized as either containing a single premise and conclusion or being a declarative statement without a conclusion. Then, we begin assessing potential issues or errors in the reasoning chains. Now, we follow an exhaustive approach to create fine-grained error categories. We begin with 30 reasoning chains (6 puzzles x 5 reasoning chains from LLMs) to manually identify potential errors. Next, we categorize these errors in a structured format. We then add another 30 reasoning chains to see if any new types of errors emerge. If new errors are identified, we refine our categories accordingly. This process is repeated until we evaluate a total of 150 reasoning chains and no new types of errors are found. Based on this method, we have carefully filtered and categorized several errors made by LLMs, presenting them as five broad categories and nine sub-categories.
@@ 4.1 Experiments
We evaluate a range of closed-source LLMs including GPT-4-Turbo, Claude-3-Opus, and Gemini-Pro, and open-source models Llama-2-13B, and Mistral-7B-Instruct on GridPuzzle in the Zero-shot-CoT setting Kojima et al. (2022). We also conducted a scaling experiment on Llama-2-70B and the results are given in the Appendix F. Our GridPuzzle dataset consists of a set of instances denoted as $\mathcal{P}={<p_{n}^{i\times j},a_{n}>}$, where $p_{n}^{i\times j}$ is $n^{th}$ puzzle instance with grid size of $i\times j$ and $a_{n}$ as a gold answer. We prompt each LLM to generate a reasoning chain before predicting answer $\hat{a}$. To evaluate each model in the Zero-shot-CoT setting, we provide $<I,p_{n}^{i\times j}>$ as input to the model and predict an answer $\hat{a}$ where $I$ is a natural language instruction. The evaluation is conducted on the OpenAI, Google, and Anthropic model versions released in April 2024 with temperature setting 0 for deterministic predictions. NVIDIA A100 GPUs are used for conducting the inference of open-source models with a batch size of 4. The example prompts used for these experiments are provided in Appendix A.
Figure 2: The process of calculating PuzzleEval metrics is described above. The reasoning chains are produced by our five LLMs and the gold solution is taken from our GridPuzzle dataset.
@@ 4.2 Metrics
Figure 3: Performance of five different LLMs in terms of accuracy on the GridPuzzle dataset.
@@ 5.2 Reasoning Chain Evaluation
Figure 4: The percentage distribution of the broad error categories across the combined reasoning steps of all five LLMs. The total number of steps generated by each model is provided inside the round brackets below the model names.
@@
The accuracy is low despite the reasoning chains being mostly error-free.
The disparity between accuracy and PuzzleEval arises from the relative location of errors within the reasoning chains. It has been observed that “RR” category reasoning steps mainly occur in the initial half of the chain, leading to a high overall PuzzleEval score. Conversely, errors in the “RW”, “WR”, and “WW” categories typically occur in the latter half, resulting in incorrect final answers and lower accuracy scores. Based on error taxonomy, “RW”, “WR”, and “WW” broad error categories have been further dissected into $6\times 3$ error sub-categories, with their distribution across reasoning steps shown in Figure 5.
Dominant broad categories: RW and WW.
The most common error sub-category across all heatmaps appears to be the “-”, the absence of errors. All the reasoning steps with “NC” and “RR” classifications fall in this category. To observe the actual overall trend across all 5 LLMs, the top 10 most common error sub-categories have been listed in the last sub-figure of Figure 5. The top categories ‘a’ and ‘c’ refer to the Wrong Reasoning and the Wrong Elimination sub-categories under the “RW” category. These errors arise when the premise is correct but the LLMs fail to make accurate deductions from it. A number of the top 10 sub-error categories (‘1a’, ‘4a’, ‘4b’, ‘4c’, ‘5a’, ‘5c’, and ‘6a’) emerge from the “WW” category.
@@
Program-of-Thought 10 -
Table 5: The results for accuracy and PuzzleEval using GPT-4-Turbo, with and without mitigation strategies for the 60 samples of $3\times 4$ grid-size.
Proprietary LLMs are way better at GridPuzzle than Open-Source LLMs.
From the results of objective and subjective metrics, it is evident that the open-source models have lower performance on the grid-puzzle-solving task than the proprietary models. The Llama-2 and Mistral models have the lowest accuracy values and their low performance on the PuzzleEval consistently degrades with the increase in the size and complexity of the grids. The Claude-3, Gemini, and GPT-4 models have higher values of accuracy but their performance across the grid sizes in the PuzzleEval is inconsistent. The disparity in the performance of both kinds of models can be attributed to the difference in their parameter sizes and the low instruction following capabilities of small open-source models.
Popular reasoning error mitigation strategies do not improve LLMs on GridPuzzle.
We conduct a case study on a subset of GridPuzzle, focusing on a 3x4 grid size, utilizing commonly employed prompting techniques to enhance the reasoning capabilities of LLMs. In particular, we use five strategies: (1) Plan-and-Solve Wang et al. (2023), (2) Self-correct Zhang et al. (2024), (3) Self-discover Zhou et al. (2024), (4) Feedback-Learning, and (5) Program of Thought prompting Chen et al. (2023). We updated the prompts corresponding to these techniques to include some of our major findings from the reasoning chain evaluations and error categorization analysis as precautionary instructions.
The first strategy Plan-and-Solve prompts the model to first generate a plan to solve the given problem and then follow those steps. The second strategy is inspired by the Self-correct method which uses a combination of self-verification and self-refine to improve reasoning. Next, we used the Self-discover technique which is a 2-step structured reasoning process. We created our prompting technique called “Feedback-Learning” by providing specific feedback system instructions to the LLM based on our error taxonomy. Lastly, we also implemented a code-style prompting technique that implements a code to solve the puzzle but does not give a reasoning chain. The detailed prompt structure is described in Appendix E and the results of these strategies are in Table 5. It is evident from the results that prompting-based strategies are not sufficient to significantly improve the LLM reasoning on the grid-puzzle-solving task. Compared to the rest of the strategies, Self-Discover marginally improves the performance on both accuracy and PuzzleEval. These results indicate the need to develop techniques beyond prompting by having deeper insights from LLMs’ reasoning chains.
@@ 6 Conclusion
The dataset, GridPuzzle, used for this study is based on 274 puzzles from the open-source platform (more details in section 3.1). No personal information from data creators has been collected during the creation of the dataset. The data collection process strictly adheres to the terms of use and privacy policies of the platform. Furthermore, the use of proprietary LLMs such as GPT-4, Gemini, and Claude-3 in this study adheres to their policies of usage. We have used AI assistants (Grammarly and ChatGPT) to address the grammatical errors and rephrase the sentences.
@@ Acknowledgement
We thank the anonymous reviewers for their constructive suggestions. We extend our gratitude to the Research Computing (RC), and Enterprise Technology at ASU for providing computing resources, and access to the ChatGPT enterprise version for experiments. We acknowledge support by a 2023 Spring Amazon Research Award (ARA). This material is also based upon work supported by the Engineering Research and Development Center - Information Technology Laboratory (ERDC-ITL) under Contract No. W912HZ24C0022.
Appendix A GridPuzzle Dataset - Sample Puzzle
@@
Appendix B Auto Evaluator: GPT-4o
@@ Appendix C Evaluation of Reasoning Chains
Appendix D Annotation Guideline
To conduct the manual analysis of the reasoning chains, the annotators were provided the guidelines described in figure 10. The same guideline was also used to create the system prompt for the GPT-4o Auto-evaluator. The annotation process was conducted by 5 annotators and the annotations were also cross-examined to resolve any discrepancies.
Figure 10: The 12-step guideline provided to the annotators to conduct manual evaluation of the reasoning chains.
Appendix E Mitigation Strategy Prompts
@@
Figure 13: The prompt structure for the Program of Thought prompting technique. In the first part, we asked the LLM to generate a code to solve the given puzzle. In the second part, an LLM is prompted to implement the code produced in the first part to get the Final Answer table.
We also implemented a code-based technique to sole GridPuzzle. We borrowed the PoT prompt from the original implementation to create a solver function to solve the puzzles. Next, we asked an LLM to implement this code and produce the Final Answer. Since the codes produced by the LLM may contain some errors we utilized the LLM’s compiler to implement the code instead of a rigid Python compiler. The prompt structure is provided in figure 13. Next is the Self-correct strategy which merges Self-verify and Self-refine qualities to minimize LLM reasoning errors. It starts with solving the puzzle using a Zero-shot-CoT prompt, followed by prompting the LLM to verify and refine the solution. Finally, it integrates the model’s suggestions with insights from our error taxonomy to enhance the puzzle-solving response. The prompt structure for this strategy is shown in figure 14. Lastly, the Self-Discover strategy, depicted in figure 15, proved most effective in reducing LLM reasoning errors in puzzle-solving. This approach begins by having the model analyze the problem and potential errors, follows with a list of prescribed reasoning modules, prompts the LLM to select and apply the most suitable module, and concludes by using a structured prompt to solve the puzzle.
@@
Appendix F Model Scaling Effect: Llama-70B
We conducted a case study on the Llama models to analyze their performance on GridPuzzle with increasing model parameter size. We repeated the same experiment in the Zeo-shot-CoT setting with the Llama-70B model. We found that the performance of the bigger model was marginally higher than the 13B model. The Accuracy went up from 1 correct final answer in the 13B model to 2 in the 70B model. The scores on PuzzleEval also went up 11% on average. However, despite the slight improvement, the Llama model’s performance was still inferior to GPT-4, Gemini, and Claude. The experimental findings are presented in Table 6. We infer that even with the increasing model parameter size, the LLMs lack the intrinsic reasoning capabilities required to solve complex logic problems such as GridPuzzle.
Model 3 x 4 3 x 5 4 x 4 4 x 5 4 x 6 Avg
Llama-70B 0.51 0.51 0.52 0.58 0.42 0.52
Table 6: The results for PuzzleEval on the different grid sizes available in GridPuzzle dataset in terms of ACS for Llama-70B. The Accuracy of Llama-70B was 2/274 puzzles.